Circular No. 6873

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998ao AND 1998ap IN ANONYMOUS GALAXIES

J. Maza, Department of Astronomy, University of Chile, reports the discovery of two supernovae by Roberto Antezana on a T-Max 400 films taken by L. E. Gonzalez with the Maksutov telescope at Cerro El Roble. SN 1998ao (B ∼ 18) was found on a film taken on Mar. 25.07 UT, and it is located at α = 10^{h}07^{m}14^{s}.58, δ = −26^{o}23′35′′.1 (equinox 2000.0), which is 9′′.7 east and 3′′.7 south of the center of the host galaxy. The object was confirmed by R. Covarrubias at Cerro Tololo using the 0.91-m telescope on Mar. 28, and it is not present (B > 20) on a film obtained on Feb. 25. A spectrum of SN 1998ao was obtained by J. Maza on Apr. 2.15 at Cerro Tololo using the Blanco 4-m telescope; M. Phillips classified the spectrum as a normal type-II supernova, younger than 10 days after maximum (the redshift of the parent galaxy is z = 0.01557). SN 1998ap (B ∼ 17) was found on a film taken on Mar. 31.385, and it is located at α = 20^{h}00^{m}35^{s}.58, δ = −52^{o}42′36′′.8 (equinox 2000.0), which is 16′′.2 west and 2′′.9 north of the center of the host galaxy. SN 1998ap was confirmed by J. Maza on Apr. 3.8 at Cerro Tololo with the Blanco 4-m telescope; a spectrum shows narrow hydrogen lines (FWHM ∼ 1000 km/s) at a redshift of z = 0.04146 superimposed on a blue featureless continuum. The object is not present on a film obtained with the same telescope on 1997 Oct. 7.

SAX J1747.0–2853 A. Bazzano and P. Ubertini, Istituto di Astrofisica Spaziale, CNR,

Rome; J. in ’t Zand and J. Heise, Space Research Organization of the Netherlands (SRON), Utrecht; D. Ricci and L. A. Antonelli, BeppoSAX Science Data Center, Rome; and M. J. S. Smith, BeppoSAX Science Operation Centre, Rome, and SRON, report on a BeppoSAX follow-up observation of SAX J1747.0–2853 (cf. IAUC 6846) with the Narrow Field Instruments on Mar. 23: “An improved position has been obtained, corresponding to α = 17^{h}47^{m}02^{s}, δ = −28^{o}52′.5 (equinox 2000.0; error radius 1′); this is 0′.5 from the previously reported position and 0′.8 from the centroid position of the 1976 transient GX +0.2–0.2 (Proctor et al. 1978, MNRAS 185, 745). The source shows strong absorption, corresponding to a hydrogen column density of ∼ 7 × 10^{22} cm^{−2} and an average flux of 5.0 × 10^{−10} erg cm^{−2} s^{−1} in the band 2–10 keV. Since Mar. 23, the source was monitored with the Wide Field Camera 1 and has shown repeated bursting activity; a persistent emission corresponding to 7 mCrab in the range 2–9 keV was detected on Apr. 3 and 4.”

1998 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green